# references.bib

### v1

@@ 4 Results
We applied various combinations of masking strategy and data augmentation, the results of which are shown in Fig. 4.
Figure 4: Linear readout accuracy quantifies representation quality after pretraining. a) Influence of masking strategy in comparison to the multi-view approaches CLAPP [illing2021local] and LPL [halvagal2023combination], that both come in two variants trained with a local learning rule or backpropagation. The remaining baselines are pure autoencoding (AE) from full image to full image in the same network as the masked methods, and directly conducting linear probing on the input images. b) Influence of data augmentation. In the left- and rightmost bar, the filled and outlined bars overlap, i.e., classification accuracy is unaffected by augmentation. c) Incomplete masking by Gaussian blurring, instead of uniform average coloring, drastically decreases representation quality. Error bars indicate the standard deviation across five randomly seeded runs.
@@ 4.1 Influence of Masking Strategy
Out of the used masking strategies, we identified the masked periphery condition as the most promising candidate for a biological model. This masking technique inspired by foveal visual perception led to the second highest classification accuracy in linear readout (mean value $67.9\pm 0.4\%$) when crop-and-resize transformations were used during pretraining, outperformed only by the random patch-wise masks ($70.2\pm 0.4\%$). Strong performance in these settings supports the assumption that solving the pretraining tasks of predicting hidden image areas leads to discovery of object representations [kong2023understanding]. In contrast, deblurring the peripheral region from the foveal filter resulted in comparably poor representations with readout accuracy not much higher than pure autoencoding. Presumably, this condition allows the network to find a local solution that does not require a globally integrated representation. To more closely investigate the effect of input information content on classification accuracy, we introduced Gaussian blurred information in the masked areas (Fig. 2, third column) of the masked periphery and random patch masks. Despite the low information content, the introduced information was sufficient to degrade performance substantially (Fig. 4c).
@@ 4.2 Decorrelation of Latent Space Neurons
To investigate the degree to which neurons in latent space develop independent tuning properties, we computed the correlation term $c$ using Eq. 2. Interestingly, both masked periphery and random patches decorrelate neural representations (Fig. 5), and more strongly than the foveal filter. This decorrelation, reminiscent of sparse coding [olshausen1996emergence], emerged without being explicitly enforced as in other approaches [olshausen1996emergence, bardes2021vicreg, zbontar2021barlow].
Figure 5: Masked image modeling implicitly decorrelates latent space neurons. Shaded regions indicate the standard deviation across five randomly seeded runs.
@@ 4.3 Influence of Data Augmentation
Applying crop-and-resize augmentation during pretraining proved necessary for representation learning in the masked periphery condition. Without this augmentation technique, readout accuracy dropped from $67.9\pm 0.4\%$ to $56.7\pm 0.5\%$ (Fig. 4b). Biologically, this augmentation can be approximated through different viewing angles, especially since no assumption about temporal co-occurrence is made: differently augmented versions of an image are presented in different epochs. Interestingly, neither the foveal filter nor the random patches were affected by ablating the crop-and-resize transformation. Maintaining the position of the patch masks across epochs alone did not reduce readout accuracy ($70.3\pm 0.9\%$), but in combination with ablation of the crop-and-resize augmentation, accuracy reduced to $62.5\pm 0.8\%$. We conclude from this that the model requires exposure to a variety of prediction tasks from each object, whether through data augmentation or through changes in mask position.
@@
Restricting the pretraining loss to the main object instead of the background, as described in section 3.5, did not affect linear probing accuracy negatively, suggesting that on-object predictions are sufficient for classification. While there was no substantial difference in readout accuracy when disregarding, as opposed to when using, loss signal from the background in the masked periphery condition (Fig. 6b), the network pretrained using random patch masks profited from disregarding the background, especially in early epochs (Fig. 6a).
Figure 6: Influence of presegmentation and masking strategy on training speed. (a) When using random patches, using presegmentation to weight the reconstruction loss accelerates training. (b) This is not the case when masking the periphery. Shaded regions indicate the standard deviation across five randomly seeded runs.
@@ 4.5 Reconstruction Quality
Beyond the network’s disciminative function, its generative capacity bears functional relevance, e.g., for counterfactual reasoning, predictions of occluded image content, and imagination [pennartz2015brain]. An interesting effect from generative model learning can be observed in the networks pretrained with uniform (average colored) and blurry masks: These networks also reconstruct in locations in which no loss was calculated. While is to some degree expected when the masked area changes shape across image presentations (random patches), this is somewhat surprising in the masked periphery condition (Fig. 2, second column). Here, also the central region was reasonably reconstructed although the pretraining loss is indifferent to the network’s predictions. This holistic percept fits well with the proposed role of visual predictions in creating a stable visual representation [crapse2008frontal, pennartz2015brain].
Across masking paradigms, reconstruction accuracy was uncorrelated with readout accuracy: while the foveal filter achieved the best reconstructions (Fig. 2), it performed poorly in linear probing (Fig. 4). Given a pretraining paradigm, however, longer training on the reconstruction task also improved readout (Fig. 6).
A third angle to analyze the generative capacity of the network stems from the fact that eye-movements are often preceded by covert attention shifts, which increase information sampling from these areas [houtkamp2003gradual]. The sampled information can then be used to improve predictions about the expected image content. To test whether the model could exploit such a mechanism for improved reconstructions, we provided the model with a second circular input sampling area (covert attention) on top of the peripheral masking while constraining the error to the foreground (section 3.5). The additionally provided information indeed led to more truthful predictions (Fig. 7), i.e., reduced prediction errors (from 2.6e-2 to 2.3e-2), which ties the approach back to earlier theories of predictions as a stabilizing mechanism [crapse2008frontal]. However, adding this sampling during pretraining, did not alter readout accuracy significantly, neither when setting the attention circle randomly nor when placing it on random positions of the main object.
Figure 7: Attention-guided predictions for stable perception. (a) Top row: ground truth. Middle row: input in the masked periphery condition with a masking ratio of 80%. Bottom row: reconstructions after pretraining. (b) Same as (a) with additionally provided information via a second, non-central input circle (covert attention). The top row depicts the input mask consisting of these two circular areas (fovea and attention). Adding covert attention improves reconstructions.
@@ 5 Conclusion
In this paper, we investigated key components of masked image modeling (MIM) through the lens of biological vision. As a similarly performant alternative to the random patch masks common in artificial representation learning [he2022masked, woo2023convnext], we identified the more biologically plausible peripheral masking (Fig. 4a). Provided with differently augmented views of the input images (Fig. 4b), this approach does not pose restrictions on the temporal sequence in which inputs are encountered [halvagal2023combination] or on knowledge about the spatial extent of an object to obtain negative samples [illing2021local]. An additional advantage of the employed approach in contrast to purely discriminative methods is learning of a generative model (Fig. 2, Fig. 7). A general requirement for high classification accuracy across all investigated masking strategies is strict intransparency of input masks (Fig. 4c), just as we sometimes do not perceive objects in our field of view while focusing on another point in space [simons1999gorillas]. From a theoretical perspective, implicit decorrelation of latent space neurons (Fig. 5), together with previous observations of implicit invariance learning [kong2023understanding2], connect the MIM framework to latent regularization SSL [bardes2021vicreg], where these objectives are explicitly enforced. This raises the question whether and how synergies could be efficiently exploited. We conclude that peripheral masking and prediction is a candidate mechanism for self-supervised learning in the brain, potentially in addition to previously proposed non-generative methods [halvagal2023combination, illing2021local]. These and other multi-view approaches [lowe2019putting, siddiqui2023blockwise] have successfully conducted self-supervised learning with more local learning rules than backpropagation, raising the question whether the same can be done with MIM. Lastly, as primates employ eye-movements in a strategic and enactive way [thompson2022learning], an interesting extension would be to learn from multiple peripherally masked glimpses in a sequence through spatiotemporally masked videos [feichtenhofer2022masked].
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreement No. 945539 (Human Brain Project SGA3 to C.P. and S.B.). We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius and especially Thomas von Osch for the helpful discussions.

### v2

@@ 4 Results
To identify the minimally necessary components for biological MIM, we applied various combinations of masking strategy and data augmentation, and plot their influence on representation quality in Fig. 4a-c.
@@ 4.1 Influence of Masking Strategy
Out of the used masking strategies, we identified the masked periphery condition as the most promising candidate for a biological model. This masking technique inspired by foveal visual perception led to the second-highest classification accuracy in linear readout (mean value $67.9\pm 0.4\%$) when crop-and-resize transformations were used during pretraining, outperformed only by the random patch masks ($70.2\pm 0.4\%$). Similarly, in the ConvNeXt-v2 architecture, peripheral masking led to a high linear readout accuracy of $71.6\pm 0.6\%$ (fine-tuning:
Figure 4: Linear readout accuracy quantifies representation quality after pretraining. a) Influence of masking strategy in comparison to the multi-view approaches CLAPP [illing2021local] and LPL [halvagal2023combination], that both come in two variants trained with a local learning rule or backpropagation. The remaining baselines are pure autoencoding (AE) from full image to full image in the same network as the masked methods, and directly conducting linear probing on the input images. b) Influence of data augmentation. In the left- and rightmost bar, the filled and outlined bars overlap, i.e., classification accuracy is unaffected by augmentation. c) Incomplete masking by Gaussian blurring, instead of uniform average coloring, drastically decreased representation quality. d) Masked image modeling implicitly decorrelates latent space neurons. Error bars and shaded regions indicate the standard deviation across five randomly seeded runs.
$90.9\pm 0.3\%$), compared to $68.2\pm 0.3\%$ (fine-tuning: $89.4\pm 0.4\%$) attained with random patch masks. Strong performance in these settings supports the assumption that solving the pretraining tasks of predicting hidden image areas leads to discovery of object representations [kong2023understanding]. In contrast, deblurring the peripheral region from the foveal filter resulted in comparably poor representations, with readout accuracy not much higher than pure autoencoding. Presumably, this condition allows the network to find a local solution that does not require global integration [kong2023understanding]. To more closely investigate the effect of input information content on classification accuracy, we replaced uniform masking with Gaussian blurring (Fig. 2, third column). Despite the low information content, the introduced information was sufficient to substantially degrade performance in both the masked periphery and random patch condition (Fig. 4c).
@@ 4.2 Decorrelation of Latent Space Neurons
To investigate the degree to which neurons in latent space develop independent tuning properties, we computed the correlation term $c$ using Eq. 2. Interestingly, both masked periphery and random patches decorrelated neural representations (Fig. 4d), and more strongly than the foveal filter. This decorrelation, reminiscent of sparse coding [olshausen1996emergence], emerged without being explicitly enforced as in other approaches [olshausen1996emergence, bardes2021vicreg, zbontar2021barlow].
@@ 4.3 Influence of Data Augmentation
Applying crop-and-resize augmentations during pretraining proved necessary for representation learning in the masked periphery condition. When leaving out this augmentation, readout accuracy dropped from $67.9\pm 0.4\%$ to $56.7\pm 0.5\%$ (Fig. 4b). Biologically, cropping and resizing can be approximately obtained by changes in viewpoint and viewing angle, especially since no assumption about temporal co-occurrence is made: differently augmented versions of an image are presented in different epochs. Interestingly, neither the foveal filter nor the random patches were affected by ablating the crop-and-resize transformation. Maintaining the position of the patch masks across epochs alone did not reduce readout accuracy ($70.3\pm 0.9\%$), but in combination with ablation of the crop-and-resize augmentation, accuracy reduced to $62.5\pm 0.8\%$. We conclude that the model requires exposure to a variety of prediction tasks from each object, whether through data augmentation or through changes in mask position.
@@
Restricting the pretraining loss to the main object instead of the background (cf. section 3.7) did not affect linear probing accuracy negatively, suggesting that on-object predictions are sufficient for classification. While there was no substantial difference in readout accuracy when disregarding, as opposed to when using, loss signal from the background in the masked periphery condition (Fig. 5b), the network pretrained using random patch masks profited from disregarding the background, especially in early epochs (Fig. 5a).
Figure 5: Influence of presegmentation and masking strategy on pretraining efficiency. (a) When using random patches, using presegmentation to weight the reconstruction loss accelerated pretraining. (b) This was not the case when restricting the reconstruction loss to the foreground object. Shaded regions indicate the standard deviation across five randomly seeded runs.
@@ 4.5 Reconstruction Quality
Beyond the network’s discriminative function, its generative capacity bears functional relevance, e.g., for counterfactual reasoning, predictions of occluded image content, and imagination [pennartz2015brain]. An interesting effect from generative model learning could be observed in the networks pretrained with uniform (average colored) and blurry masks: These networks also reconstructed in locations in which no loss was calculated. While this would be expected to some degree when the masked area varies across image presentations (random patches), it is surprising in the masked periphery condition (Fig. 2, second column). Here, also the central region was reconstructed reasonably well, although it did not contribute to the pretraining loss. This holistic percept fits well with proposals on the role of visual predictions in creating a stable visual representation [crapse2008frontal, pennartz2015brain].
Across masking paradigms, reconstruction accuracy was uncorrelated with readout accuracy: while the foveal filter achieved the best reconstructions (Fig. 2), it performed poorly in linear probing (Fig. 4a). Given a pretraining paradigm, however, longer training on the reconstruction task also improved readout (Fig. 5).
A third angle to analyze the generative capacity of the network stems from the fact that eye movements are often preceded by covert attention shifts, which increase information sampling from these areas [houtkamp2003gradual]. The sampled information can then be used to improve predictions about the expected image content. To test whether the model could exploit such a mechanism for improved reconstructions, we provided the model with a second circular input sampling area (covert attention) on top of the peripheral masking while constraining the error to the foreground (section 3.7). The additionally provided information indeed led to more truthful predictions (Fig. 6), i.e., reduced prediction errors (from 2.6e-2 to 2.3e-2), which ties the approach back to earlier theories of predictions as a stabilizing mechanism [crapse2008frontal]. Applying this twofold sampling during pretraining, however, did not alter readout accuracy significantly, irrespective of whether the covert attention circle was placed at random or on the main object.
Figure 6: Sampling via covert attention improves predictions and stabilizes perception. (a) Top row: ground truth. Middle row: input in the masked periphery condition with a masking ratio of 80%. Bottom row: reconstructions after pretraining. (b) Same as (a) with additionally provided information via a second, non-central input circle (covert attention). The top row depicts the input mask consisting of these two circular areas (fovea and attention).
@@ 5 Conclusion
In this paper, we investigated key components of masked image modeling (MIM) through the lens of biological vision. As a similarly performant alternative to the random patch masks common in artificial representation learning [he2022masked, woo2023convnext], we identified the more biologically plausible peripheral masking (Fig. 4a). Provided with differently augmented views of the input images (Fig. 4b), this approach does not pose restrictions on the temporal sequence in which inputs are encountered, as does [halvagal2023combination], or on knowledge about the spatial extent of an object to obtain negative samples, as does [illing2021local]. We conclude that peripheral masking and prediction is a candidate mechanism for self-supervised learning in the brain, with saccadic prediction errors computed in the early (but likely not primary) visual cortex [ehinger2015predictions] where bottom-up signals are compared to feedback from the frontal eye fields [crapse2008frontal]. However, further experimental studies are required to definitely link these error signals to plasticity in the ventral stream, and to elucidate whether covert attention shifts can elicit similar prediction errors. Functionally, an additional advantage of the employed approach in contrast to purely discriminative methods is the acquisition of a generative model (Fig. 2, Fig. 6). That multi-view SSL has been implemented with more local learning rules than backpropagation [illing2021local, halvagal2023combination, lowe2019putting, siddiqui2023blockwise] prompts the question of whether the same can be done with MIM. From a theoretical perspective, implicit decorrelation of latent space neurons (Fig. 4d), together with previous observations of implicit invariance learning [kong2023understanding2], connect the MIM framework to SSL via latent regularization [bardes2021vicreg], where these objectives are explicitly enforced. This raises the question of whether and how synergies could be efficiently exploited in models combining both learning paradigms [assran2023self, jing2022masked]. As a general requirement for high classification accuracy across all investigated masking strategies, we identified opacity of the input masks (Fig. 4c), just as humans sometimes do not perceive objects in their field of view when attending to another point in space [simons1999gorillas]. As primates employ eye movements in a strategic and enactive way [thompson2022learning], an interesting extension would be to learn from multiple peripherally masked glimpses in a sequence through spatiotemporally masked videos [feichtenhofer2022masked].
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreement No. 945539 (Human Brain Project SGA3 to C.P. and S.B.). We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius, especially Thomas von Osch for the discussions.